In the analysis of 0.5 g sample of feldspar, a mixture of chlorides of sodium and potassium is obtained, which weighs 0.1180 g. Subsequent treatment of the mixed chlorides with silver nitrate gives 0.2451 g silver chloride. What is the percentage of sodium oxide and potassium oxide in the sample?

Grade:10

1 Answers

Aditi Chauhan
9 years ago
Sol. Molarity = (Moles of solute)/(Volume of solution (L)) moles of urea = 120/60 = 2 Weight of solution = weight of solvent + weight of solute = 1000 + 120 = 1120g ⇒ Volume = 1120g/(1.15g / mL) × 1/(1000mL/L) = 0.974L ⇒ Molarity = 2.000/0.974 = 2.05M